# Zhonghong Kuang Renmin U Yi Liu

### v1

@@ Abstract
We study the optimal design of relational contracts that incentivize an expert to share specialized knowledge with a novice. While the expert fears that a more knowledgeable novice may later erode his future rents, a third-party principal is willing to allocate her resources to facilitate knowledge transfer. In the unique profit-maximizing contract between the principal and the expert, the expert is asked to train the novice as much as possible, for free, in the initial period; knowledge transfers then proceed gradually and perpetually, with the principal always compensating the expert for his future losses immediately upon verifying the training he provided; even in the long run, a complete knowledge transfer might not be attainable. We further extend our analysis to an overlapping-generation model, accounting for the retirement of experts and the career progression of novices.
@@ 1 Introduction
To shed light on these questions, we study a stylized model where a principal (she) interacts over time with two agents, an expert (he) and a novice (he). The expert maintains the maximum knowledge level throughout the game, and the novice begins with a knowledge level of 00. In each period, the principal makes a payment to the expert, and then the expert trains the novice, which increases the novice’s knowledge in the next period to a level at the expert’s choice. The stage payoffs depend on the novice’s current knowledge level: a more knowledgeable novice leads to higher stage payoffs for both the novice and the principal, but a lower stage payoff for the expert, capturing the potential loss of his task assignment/market share to a more capable novice. In this case, the expert is willing to train the novice only if he will be compensated for the resulting loss of future stage payoffs. We also assume that the bilateral surplus between the principal and the expert increases with the novice’s knowledge level. [^3] In the baseline model, the novice is a passive player with no decisions to make. In Appendix A, we provide several microfoundations where the novice becomes an active player.
@@
We model knowledge transfer as an irreversible and non-contractible process with delayed observation: After some training is provided in a period, it is not observed until the next period, and due to its almost impossibility to verify in court, the parties cannot write an enforceable contract that depends on the novice’s knowledge level. Consequently, there is a lack of commitment on both sides, making immediate knowledge transfer impossible. [^4] If knowledge were fully transferred to the novice immediately, the principal would have no incentives to compensate the expert afterwards. It is also impossible to ask the principal to pay in advance because the expert, expecting no further payments, would not have incentives to train the novice. This dual commitment failure necessitates gradual knowledge transfer through self-enforcing relational contracts between the principal and the expert, restricted by their respective incentive constraints. Specifically, the expert is willing to train the novice for future payments; meanwhile, the principal prefers to pay the expert to continue training rather than terminating both payments and training.
@@
We fully characterize the unique profit-maximizing contract (Theorem 1). In the optimal contract, knowledge is gradually transferred through an infinite-period arrangement. [^5] Except when the players are highly impatient, in which case no knowledge transfer can be incentivized. In the first period, the principal asks the expert to transfer as much knowledge as possible, for free, to the novice. After that, in each period, the principal compensates the expert for the present value of his future payoff losses incurred from the last period’s knowledge transfer; meanwhile, the principal requests the expert to transfer additional knowledge, the present value of which to the principal exactly equals the compensation provided. The optimal knowledge sequence dominates, period by period, all other knowledge sequences that can be implemented by some relational contract, and because of that, it also serves as the unique knowledge sequence in every Pareto-efficient contract (Theorem 2).
@@
To understand these properties, note first that no payment is made in the very first period because it does not affect the expert’s incentives to train in any period. [^6] This result is reminiscent of the “knowledge gift” in Garicano and Rayo (2017), but they occur for very different reasons. See related literature and Section 5.2 for a discussion. After the initial period, the principal always fully compensates the expert for the present value of his future losses, immediately after the training he provided is verified. In principle, such a compensation could take multiple periods, but doing so would only strengthen the principal’s incentives to default in the future, so it is desirable to frontload. Moreover, through a perturbation argument, we show that any optimal contract must bind the expert’s incentive constraint in every period, as well as the principal’s incentive constraints starting from the second period. These binding constraints pin down a difference equation for knowledge levels in adjacent periods, which allows us to trace out a sequence of knowledge levels from any initial condition. Finally, it turns out that a greater amount of training in the initial period will result in higher knowledge levels in all future periods, so that the principal’s optimization problem boils down to incentivizing the highest possible amount of training at the beginning of the game.
@@
The qualitative properties of the optimal contract align closely with the quid pro quo (“technology for market access”) policy adopted in the Chinese automobile industry over the last four decades. A prominent example is the joint venture established in 1997 between Shanghai Automobile Industry Corporation (SAIC) and General Motors (GM), which led to the creation of the SAIC-GM brand for manufacturing passenger vehicles in China. Consistent with our theoretical predictions, the technology transfer process unfolded gradually over many years; meanwhile, the local government provided continuous support to GM, primarily through land price discounts, tax incentives, and preferential procurement treatment. This arrangement ultimately enabled SAIC to launch its own independent automotive brand, Roewe, in 2006, although cars under this local brand were generally regarded as lower in quality compared to those under the SAIC-GM brand. In fact, even by 2012, GM had not yet transferred some of its key technologies. [^7] As reported, “GM is holding tight to its more valuable technology. Beijing is eager to tap into foreign auto companies’ clean-energy technologies, [b]ut GM doesn’t want to share all its research with its Chinese partner.” See Sharon Terlep, “Balancing the Give and Take in GM’s Chinese Partnership.” Wall Street Journal, August 19, 2012. The original joint venture agreement also specifically required GM to offer know-how and product lines as equity, thereby pushing for a significant amount of technology transfer at the initial stage of the partnership. We discuss this application in more detail in Section 3.4, and refer the reader to Bai, Barwick, Cao and Li (2025) for a thorough explanation of the institutional background.
@@
We characterize the profit-maximizing stationary contract and show that it maintains most of the qualitative features from the baseline model, with two substantial differences (Theorem 3). First, the expert fully trains the novice by the time he retires. This happens because a knowledgeable novice cannot hurt the expert anymore after his retirement. Second, the equilibrium’s qualitative properties no longer depend on the discount factor, and in particular, the trivial contract without any transfers is always suboptimal, even for highly impatient players. Although the principal never incurs the catch-up cost in the optimal contract, the presence of such a cost serves as a credible punishment on the principal’s default, which helps sustain positive knowledge transfers throughout the relationship. [^8] If there were no such costs, there would only be positive knowledge transfer in the final period of the expert phase. Moreover, we find that uniformly increasing the catch-up cost function will benefit the principal and speed up the knowledge transfer process (Proposition 2).
@@
Our paper differs from these studies in three key respects. First, in the aforementioned papers, the novice is the sole productive player whose productivity increases with his knowledge level, so the expert has an intrinsic motivation to train him. In contrast, we highlight a fundamental conflict of interest between the two, arising from the fact that a more capable novice may become a stronger competitor to the expert. Yet, it is in the interest of a third-party principal to promote knowledge transfer. Second, in those models, the expert serves as the contract designer who uses knowledge transfers as a main source of reward for the novice’s production. In our model, however, the contract designer is the principal who leverages monetary transfers to incentivize the expert to train the novice. Third, Garicano and Rayo (2017) assume that the expert has one-period commitment, and Fudenberg and Rayo (2019) and Fudenberg, Georgiadis and Rayo (2021) endow the expert with full commitment power. By contrast, our model assumes no commitment on either side, reflecting the non-contractible nature of knowledge. A more detailed comparison between our results and theirs is provided in Section 5.2.
@@ 2.1 Players, Actions, and Payoffs
A principal (she) interacts with an expert (he) and a novice (he) over infinite periods $t=0,1,\cdots$, and they discount future payoffs with a common discount factor $\delta\in(0,1)$. The expert possesses one unit of perfectly-divisible productive knowledge, and he can train the novice to increase the novice’s knowledge level at any rate he desires. The novice’s initial knowledge level at time 00 is $s_{0}=0$. Let $s_{t}\in[0,1]$ be the novice’s knowledge level at the beginning of period $t$, and let $x_{t}\equiv s_{t+1}-s_{t}$ be the amount of knowledge accrued by the novice in period $t$. The expert’s knowledge level remains at $1$ in every period regardless of the training he provides.
@@
$s_{t}$ is observedExpert gets $w(s_{t})+p_{t}$rincipal gets $\pi(s_{t})-p_{t}$rincipal chooses $p_{t}$xpert chooses $x_{t}$$s_{t+1}$ is observed($x_{t}$ is observed)
@@
Because knowledge transfer is difficult to observe immediately, and even more so to contract on, we assume that the principal does not observe $x_{t}$ until the beginning of $t+1$ and the players are unable to write an enforceable long-term contract based on any $x_{t}$. Instead, in period 00, the principal and the expert agree on a relational contract (or simply, a contract): a self-enforcing agreement that specifies, for each period, a knowledge stock $s_{t}$ and a compensation $p_{t}$, conditional on the players staying on path. Let $\bm{s}=(s_{0},s_{1},\cdots)$ and $\bm{p}=(p_{0},p_{1},\cdots)$, and we denote a contract by $(\bm{s},\bm{p})$.
@@
The novice in our model is a passive player with no choices to make, and will thus be absent from the analysis. We introduce and make assumptions on his stage payoff, $v$, only for completeness. In the microfoundations in Appendix A, the novice’s active decisions will give rise to such an $v$ and drive other assumptions on $\pi$ and $w$.
@@ 2.3 Principal’s Problem
Notice that the trivial contract $\{s_{t}=s_{0},\ p_{t}=0\}_{t\geq 0}$ is implementable, in which case the principal’s time-00 profit is $\frac{\pi(s_{0})}{1-\delta}$. The principal aims to select an implementable contract that maximizes her time-00 profit, which leads to the following program:
@@ 3.1 Necessary Conditions
If training occurs at any point, it must never fully stop. Suppose instead that $x_{T-1}>0$ but no further training is provided starting from period $T$. To be willing to provide $x_{T-1}$, the expert has to be compensated for the loss of his future stage payoffs through $\{p_{t}\}_{t\geq T}$. However, because $x_{t}=0$ for all $t\geq T$, the principal would find it optimal not to make any payment from period $T$ onward, rendering the training at $T-1$ impossible. This observation implies that the first-best outcome (i.e., $s_{t}=1,\ \forall t\geq 1$) is not implementable. Even worse, knowledge cannot be fully transferred in finite time.
@@
Next, we examine the players’ incentive constraints in detail. Take any implementable $(\bm{s},\bm{p})$, and let us assume that (E-IC) is slack in some period $k$. We now show that the principal’s profit can be improved in two steps.
Step 1. Fixing $\bm{s}$, consider the following payment scheme $\bm{p}^{*}(\bm{s})$:
@@
[…]
Step 2. Suppose for simplicity that $s_{k+1}<s_{k+2}$. [^10] The proof in Appendix B.1 deals with the general case where $s_{k+1}\leq s_{k+2}$. Consider another perturbation $(\bm{s}^{\prime},\ \bm{p}^{*}(\bm{s}^{\prime}))$ which only increases $s_{k+1}$ by a small amount ($s_{k+1}^{\prime}=s_{k+1}+\varepsilon$; $s_{t}^{\prime}=s_{t}$ if $t\neq k+1$). Note that this perturbation improves the principal’s profit by increasing bilateral surplus while keeping the expert’s time-00 value unchanged. [^11] Indeed, whenever the payment scheme (P*) is used, all (E-IC) constraints are binding, implying that $W_{0}=w(s_{0})+\delta W_{1}=\frac{w(s_{0})}{1-\delta}$. We now explain how it still maintains all the (P-IC) constraints. Because (P-IC) was made slack at $k+1$ by the first step, it continues to hold in that period despite a small perturbation. In addition, other things equal, a higher $s_{k+1}$ implies a lower $x_{k+1}$, so that the compensation to be made in period $k+2$ becomes lower according to (P*), relaxing the principal’s incentive constraint at $k+2$. In every period prior to $k+1$, the principal’s incentive constraint is also relaxed, because future surplus gets higher while the expert’s continuation value remains the same as before. Finally, the (P-IC) constraints after $k+2$ are unaffected by the prior changes.
Taken together, this two-step perturbation delivers an implementable, profit-improving contract whenever some of the (E-IC) or (P-IC) constraints are slack, hence the claim.
@@
Economically, this payment scheme frontloads the expert’s compensation by paying him the present value of all his future losses immediately after the outcome of his previous training is observed (i.e., $x_{t-1}\equiv s_{t}-s_{t-1}$). To see why this is optimal, note that for each knowledge transfer the expert provides, he has to be compensated for his loss, $w(s_{t-1})-w(s_{t})$ per period, through discounted future payments, for otherwise he can simply stop training the novice. If the compensation for $x_{t-1}$ were to be made in multiple payments, say, at $t$ and $t+1$, this would imply that the expert’s incentive constraint at $t$ must be slack, because now the expert’s deviation in period $t$ would make him forfeit, at $t+1$, not only the compensation for $x_{t}$ but also part of the compensation for $x_{t-1}$. Consequently, the principal can frontload some payment from period $t+1$ to $t$, which would relax her (P-IC) at $t+1$ without violating any other constraints. The relaxation of the principal’s incentive constraint then creates room for a strict improvement of the principal’s profit (via Step 2 below Claim 2).
@@
To understand (BE), recall that the principal needs to pay the expert $\frac{w(s_{t-1})-w(s_{t})}{1-\delta}$ in period $t$. Only after this payment is made, the expert will carry out the training of $x_{t}$, which increases the novice’s knowledge from $s_{t}$ to $s_{t+1}$. From the principal’s perspective, this training generates a perpetual benefit of $\pi(s_{t+1})-\pi(s_{t})$ each period from $t+1$ onward, which totals $\frac{\delta[\pi(s_{t+1})-\pi(s_{t})]}{1-\delta}$ from period $t$’s standpoint. In any optimal contract, condition (BE) requires that, at every $t\geq 1$, the principal’s benefit from the expert’s training must equal the payment she is supposed to make in that period.
@@
Intuitively, the frontloading payment scheme $\bm{p}^{*}(\bm{s})$ anchors the expert’s time-00 value to $\frac{w(s_{0})}{1-\delta}$, so the principal only needs to look for an implementable contract that maximizes surplus. Once $s_{1}$ is fixed, the difference equation given by (BE) fully pins down the entire knowledge sequence. It turns out that a greater $s_{1}$ will lead to a uniformly higher sequence; therefore, it is sufficient to boost $s_{1}$ (or equivalently, the training in period 00) as much as possible.
Since $p_{0}^{*}=0$, the training in the initial period is provided “for free,” or in the language of Garicano and Rayo (2017), the expert offers a “knowledge gift.” Lemma 5 indicates that to maximize profit, the principal should induce the expert to provide as much knowledge gift as possible, so long as the contract remains implementable. [^13] We discuss how this prediction is reflected in practice in Section 3.4 and compare it to Garicano and Rayo (2017) in Section 5.2.

### v2

@@ Abstract
We study the optimal design of relational contracts that incentivize an expert to share specialized knowledge with a novice. While the expert fears that a more knowledgeable novice may later erode his future rents, a third-party principal is willing to allocate her resources to facilitate knowledge transfer. In the unique profit-maximizing contract between the principal and the expert, the expert is asked to train the novice as much as possible, for free, in the initial period; knowledge transfers then proceed gradually and perpetually, while the principal offers lump-sum compensations to the expert right after verifying each transfer; even in the long run, a complete knowledge transfer might not be attainable. Our analysis sheds light on the success of several prominent cross-border technology transfers that took place in China’s auto industry and Korea’s high-speed rail development.
@@ 1 Introduction
To shed light on these questions, we study a stylized model where a principal (she) interacts over time with two agents, an expert (he) and a novice (he). The expert maintains the maximum knowledge level throughout the game, and the novice begins with a knowledge level of 0. In each period, the principal makes a payment to the expert, and then the expert trains the novice, which increases the novice’s knowledge in the next period to a level at the expert’s choice. The stage payoffs depend on the novice’s current knowledge level: a more knowledgeable novice leads to higher stage payoffs for both the novice and the principal, but a lower stage payoff for the expert, capturing the potential loss of his task assignment/market share to a more capable novice. In this case, the expert is willing to train the novice only if he will be compensated for the resulting loss of future stage payoffs. We also assume that the bilateral surplus between the principal and the expert increases with the novice’s knowledge level. [^3] In the baseline model, the novice is a passive player with no decisions to make. In Appendix A, we provide several microfoundations where the novice becomes an active player.
@@
We model knowledge transfer as an irreversible and non-contractible process with delayed observation: After the novice is trained in a period, the outcome is not observed until the next period; moreover, because knowledge levels and increments are difficult to prove in court, the parties cannot write an enforceable contract that depends on the novice’s knowledge level. Consequently, there is a lack of commitment on both sides, making immediate knowledge transfer impossible. [^4] If knowledge were fully transferred to the novice immediately, the principal would have no incentives to compensate the expert afterwards. It is also impossible to ask the principal to pay in advance because the expert, expecting no further payments, would not have incentives to train the novice. This dual commitment failure necessitates gradual knowledge transfer through self-enforcing relational contracts between the principal and the expert, restricted by their respective incentive constraints. Specifically, the expert is willing to train the novice for future payments; meanwhile, the principal prefers to pay the expert in order to continue the transfer process rather than terminating both payments and training.
@@
We fully characterize the unique profit-maximizing contract (Theorem 1). In the optimal contract, knowledge is gradually transferred through an infinite-period arrangement. [^5] Except when the players are highly impatient, in which case no knowledge transfer can be incentivized. In the first period, the principal asks the expert to transfer as much knowledge as possible to the novice, without paying him right away. After that, in each period, the principal compensates the expert for the present value of his payoff losses incurred from the last period’s knowledge transfer; meanwhile, the principal requests the expert to transfer additional knowledge, the present value of which to the principal exactly equals the compensation provided. The optimal knowledge sequence dominates, period by period, all other knowledge sequences that can be implemented by some relational contract, and because of that, it also serves as the unique knowledge sequence in every Pareto-efficient contract (Theorem 2).
@@
To understand these properties, note that the principal pays nothing in the very first period because it does not affect the expert’s subsequent incentives to train. [^6] This result is reminiscent of the “knowledge gift” in Garicano and Rayo (2017), but they occur for very different reasons. See related literature and Section 6.1 for a discussion. After the initial period, the principal always fully compensates the expert for the present value of his future losses, immediately after the training he provided is verified. In principle, such a compensation could take multiple periods, but doing so would only strengthen the principal’s incentives to default in the future, so it is desirable to frontload. Moreover, through a perturbation argument, we show that any optimal contract must bind the expert’s incentive constraint in every period, as well as the principal’s incentive constraints starting from the second period. These binding constraints pin down a difference equation for knowledge levels in adjacent periods, which allows us to trace out a sequence of knowledge levels from any initial condition. Finally, it turns out that a greater amount of training in the initial period will result in higher knowledge levels in all future periods, so that the principal’s optimization problem boils down to incentivizing the highest possible amount of training at the beginning of the game.
@@
The qualitative properties of the optimal contract align closely with the quid pro quo (“technology for market access”) policy adopted in the Chinese automobile industry over the last four decades. A prominent example is the joint venture established in 1997 between Shanghai Automobile Industry Corporation (SAIC) and General Motors (GM), which led to the creation of the SAIC-GM brand for manufacturing passenger vehicles in China. Consistent with our theoretical predictions, the technology transfer process unfolded gradually over many years; meanwhile, the local government provided continuous support to GM, primarily through land price discounts, tax incentives, and preferential procurement treatment. This arrangement ultimately enabled SAIC to launch its own independent automotive brand, Roewe, in 2006, although cars under this local brand were generally regarded as lower in quality compared to those under the SAIC-GM brand. In fact, even by 2012, GM had not yet transferred some of its key technologies. The original joint venture agreement also specifically required GM to offer know-how and product lines as equity, thereby pushing for a significant amount of technology transfer at the initial stage of the partnership. Another notable example is the development of Korea Train Express, South Korea’s high-speed rail system. The Korean government required transfer of technology in its agreement with Alstom, and the implementation process also seems to match well with our predictions. We discuss these applications in detail in Section 4.
@@
We characterize the profit-maximizing stationary contract and show that it preserves most of the qualitative features from the baseline model, with two substantial differences (Theorem 3). First, the expert fully trains the novice by the time he retires. This happens because a knowledgeable novice cannot hurt the expert anymore after his retirement. Second, the equilibrium’s qualitative properties no longer depend on the discount factor, and in particular, the trivial contract without any transfers is always suboptimal, even for highly impatient players. Although the principal never incurs the catch-up cost in the optimal contract, the presence of such a cost serves as a credible punishment on the principal’s default, which helps sustain positive knowledge transfers throughout the relationship. [^7] If there were no such costs, there would only be positive knowledge transfer in the final period of the expert phase. Moreover, we find that uniformly increasing the catch-up cost function will benefit the principal and speed up the knowledge transfer process (Proposition 2). An application to craftsmanship preservation by charitable foundations is discussed.
@@
Our paper differs from these studies in three key respects. First, in the aforementioned papers, the novice is the sole productive player whose productivity increases with his knowledge level, so the expert has an intrinsic motivation to train him. In contrast, we highlight a fundamental conflict of interest between the two, arising from the fact that a more capable novice may become a stronger competitor to the expert. Yet, in our setting, it is in the interest of a third-party principal to promote knowledge transfer. Second, in those models, the expert serves as the contract designer who uses knowledge transfers as a main source of reward for the novice’s production. In our model, however, the contract designer is the principal who leverages monetary transfers to incentivize the expert to train the novice. Third, Garicano and Rayo (2017) assume that the expert has one-period commitment, and Fudenberg and Rayo (2019) and Fudenberg, Georgiadis and Rayo (2021) endow the expert with full commitment power. By contrast, our model assumes no commitment on either side, reflecting the non-contractible nature of knowledge. A more detailed comparison between our results and theirs is provided in Section 6.1.
@@ 2.1 Players, Actions, and Payoffs
A principal (she) interacts with an expert (he) and a novice (he) over infinite periods $t=0,1,\cdots$, and they discount future payoffs with a common discount factor $\delta\in(0,1)$. The expert possesses one unit of perfectly-divisible productive knowledge, and he can train the novice to increase the novice’s knowledge level at any rate he desires. The novice’s initial knowledge level at time 0 is $s_{0}=0$. Let $s_{t}\in[0,1]$ be the novice’s knowledge level at the beginning of period $t$, and let $x_{t}\equiv s_{t+1}-s_{t}$ be the amount of knowledge accrued by the novice in period $t$. The expert’s knowledge level remains at $1$ in every period regardless of the training he provides.
@@
$s_{t}$ is observedExpert gets $w(s_{t})+p_{t}$Principal gets $\pi(s_{t})-p_{t}$Principal chooses $p_{t}$Expert chooses $x_{t}$$s_{t+1}$ is observed($x_{t}$ is observed)
@@
Because knowledge transfer is difficult to observe immediately, and even more so to contract on, we assume that the principal does not observe $x_{t}$ until the beginning of $t+1$ and that the players are unable to write an enforceable long-term contract based on any $x_{t}$. Instead, in period 0, the principal and the expert agree on a relational contract (or simply, a contract): a self-enforcing agreement that specifies, for each period, a knowledge stock $s_{t}$ and a compensation $p_{t}$, conditional on the players staying on path. Let $\bm{s}=(s_{0},s_{1},\cdots)$ and $\bm{p}=(p_{0},p_{1},\cdots)$, and we denote a contract by $(\bm{s},\bm{p})$.
@@
The novice in our model is a passive player with no choices to make, and will thus be absent from the analysis. We introduce and make assumptions on his stage payoff, $v$, only for completeness. In the microfoundations in Appendix A, the novice’s active decisions will give rise to such an $v$ and also drive the assumptions on $\pi$ and $w$.
@@ 2.3 Principal’s Problem
Notice that the trivial contract $\{s_{t}=s_{0},\ p_{t}=0\}_{t\geq 0}$ is implementable, in which case the principal’s time-0 profit is $\frac{\pi(s_{0})}{1-\delta}$. The principal aims to select an implementable contract that maximizes her time-0 profit, which leads to the following program:
@@ 3.1 Necessary Conditions
If training occurs at any point, it must never fully stop. Suppose instead that $x_{T-1}>0$ but no further training is provided starting from period $T$. To be willing to provide $x_{T-1}$, the expert has to be compensated for the loss of his future stage payoffs through $\{p_{t}\}_{t\geq T}$. However, because $x_{t}=0$ for all $t\geq T$, the principal would find it optimal not to make any payment from period $T$ onward, rendering the training at T1T-1 impossible. This observation implies that the first-best outcome (i.e., $s_{t}=1,\ \forall t\geq 1$) is not implementable. Even worse, knowledge cannot be fully transferred in finite time.
@@
Next, if we begin with an implementable $(\bm{s},\bm{p})$ that leaves some of the (E-IC) or (P-IC) constraints slack, as a first step we can always adjust the payment scheme to $\bm{p}^{*}(\bm{s})$, defined as follows:
@@
In each period $t$, the expert receives a lump-sum payment equal to the present value of all his payoff losses that arise from the training he provided in the previous period. By construction, the contract $(\bm{s},\ \bm{p}^{*}(\bm{s}))$ makes all the (E-IC) constraints binding. Because some of the incentive constraints were slack and the discounted total surplus is kept unchanged, the shift to $(\bm{s},\ \bm{p}^{*}(\bm{s}))$ ensures that (P-IC) must be slack in some period $k\geq 1$.
Now consider increasing $s_{k}$ by a small amount and adjusting the payment scheme accordingly based on (P*). This perturbation does not violate (P-IC) in period $k$ because it was made slack by the previous step, and it only relaxes the (P-IC) constraints before and after kk. [^8] Indeed, other things equal, a higher $s_{k}$ implies a lower $x_{k}$, so the compensation to be made in period $k+1$ becomes lower according to (P*); this relaxes the principal’s incentive constraint at $k+1$. In every period prior to $k$, the principal’s incentive constraint is also relaxed because her continuation value becomes higher (future surplus gets higher while the expert’s continuation value is kept unchanged). Taken together, whenever some of the (E-IC) or (P-IC) constraints are slack, this two-step adjustment will deliver an implementable, profit-improving contract, hence the claim.
@@
Economically, this payment scheme frontloads the expert’s compensation by offering a lump-sum payment for the present value of all his future losses, right after the outcome of his previous training is observed (i.e., $x_{t-1}\equiv s_{t}-s_{t-1}$). To see why this is optimal, note that for each knowledge transfer the expert provides, he has to be compensated for his loss, $w(s_{t-1})-w(s_{t})$ per period, through a discounted sum of future payments. If the compensation for $x_{t-1}$ were to be made in multiple installments, say, at $t$ and $t+1$, this would imply that the expert’s incentive constraint at $t$ must be slack, because now the expert’s deviation in period $t$ would make him forfeit, at $t+1$, not only the compensation for $x_{t}$ but also part of the compensation for $x_{t-1}$. Consequently, the principal can frontload some payment from period $t+1$ to $t$, which would relax her (P-IC) at $t+1$ without violating any other constraints. The relaxation of the principal’s incentive constraint then creates room for a strict improvement of the principal’s profit (via the perturbation below Claim 2).
@@
To understand (BE), recall that the principal needs to pay the expert $\frac{w(s_{t-1})-w(s_{t})}{1-\delta}$ in period $t$. Only after this payment is made, the expert will carry out the training of $x_{t}$, which increases the novice’s knowledge from $s_{t}$ to $s_{t+1}$. From the principal’s perspective, this training generates a perpetual benefit of $\pi(s_{t+1})-\pi(s_{t})$ each period from $t+1$ onward, which totals $\frac{\delta[\pi(s_{t+1})-\pi(s_{t})]}{1-\delta}$. In any optimal contract, condition (BE) requires that, at every $t\geq 1$, the principal’s benefit from the expert’s training must equal the payment she is supposed to make in that period.
@@
Lemma 5 suggests that the principal should push for substantial knowledge transfer in the initial stage, so long as the scheme remains sustainable. Recall that $p_{0}^{*}=0$, so the initial transfer is provided “for free,” or in the language of Garicano and Rayo (2017), the expert offers a “knowledge gift.” A profit-maximizing principal should then try to induce the expert to deliver the largest possible knowledge gift. [^10] We discuss how this prediction is reflected in practice in Section 4 and compare it to Garicano and Rayo (2017) in Section 6.1.
Technically, the frontloading payment scheme $\bm{p}^{*}(\bm{s})$ anchors the expert’s time-0 value to $\frac{w(s_{0})}{1-\delta}$, so the principal only needs to look for an implementable contract that maximizes surplus. Once $s_{1}$ is fixed, the difference equation given by (BE) fully pins down the entire knowledge sequence. It turns out that a greater $s_{1}$ will lead to a uniformly higher sequence; therefore, it is sufficient to boost $s_{1}$ (or equivalently, the training in period 0) as much as possible.